OpenStudy (anonymous):

Brandon, Rudy, Wyatt are playing baseball. Wyatt is standing at third base. He is 43 feet from Brandon who is in center field, and 75 feet from Rudy who is at first base. How far away is Brandon from Rudy if the angle where Wyatt is standing is 65 degree?

OpenStudy (anonymous):

 See the picture. We want the length of the line between Brandon and Rudy (x) and we know the other two sides. So we know, via Pythagoras, that: 43^2 + x^2 = 75^2. You might wonder why it looks like this. Well, the 75-ft side must be the hypotenuse because it's adjacent to a less-than-90 degree angle. So the equation has 75^2 on one side. Now, the equation: 1849 + x^2 = 5625 x^2 = 5625 - 1849 = 3776 x = 61.45

OpenStudy (anonymous):